OpenStudy (anonymous):

how to solve lim x approch to 0 where function is (1+2x) ki power 1/x

OpenStudy (anonymous):

what is ki?

OpenStudy (anonymous):

1/x is the power of the function (1+2x)

OpenStudy (anonymous):

pls fast

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}(2x+1)^\frac{ 1 }{ x }\] We know that the \[\lim_{x \rightarrow 0}(1+x)^\frac{ 1 }{ x }=e\] Then we can rewrite the first limit to make it look like the second one: \[\lim_{x \rightarrow 0}(1+2x)\frac{ 2 }{ 2x }=e^2\]

OpenStudy (anonymous):

Sorry, (2/2x) is the power of (1+2x)

OpenStudy (anonymous):

thank you buddy

OpenStudy (anonymous):

bro pls tell me Limit x tends to 0 where function is  how to solve